hi can you help me solve this twp step equation: 3a-4=-16
Answer by mananth(16946)

3a-4=-16
add 4
3a-4+4=-16+4
3a=-12
divide by 3
3a/3 = -12/4
a= -4
...
m.ananth@hotmail.c